OpenStudy (anonymous):

find the integral from x to x^2 (1/t)*ln((1-t)/32)dt

OpenStudy (anonymous):

find the\[\int\limits\limits_{x}^{x ^{2}} [1/t]*\ln [(1-t)/32]\]

OpenStudy (anonymous):

d/dx of ln(1-t)=-1/(1-t) and not -1/t

OpenStudy (anonymous):

So true. Sorry, thinking

OpenStudy (anonymous):

hey i by the way i have to differentiate that whole thing

OpenStudy (anonymous):

\[d/dx(\int\limits_{x}^{x^2}[1/t]*[\ln ((1-t)/32)])\]

OpenStudy (anonymous):

could you help now

OpenStudy (anonymous):

That is easier, as they cancel out. We just have to be a bit careful.

OpenStudy (anonymous):

but the limit is from x to x^2 what about that where will the limits go

OpenStudy (anonymous):

Fundamental Theorem of Calculus, I call you to help my battle http://www.youtube.com/watch?v=PGmVvIglZx8

OpenStudy (anonymous):

i know that integral and differential will get canceled but what will t change to will it be x or x^2 or x-x^2 or something else thats what i want to know

OpenStudy (anonymous):

thamks for that link

OpenStudy (anonymous):

I am just watching that link, it has been fairly long time since I did problems like this.

OpenStudy (anonymous):

OK this video was really handy. I hope you can do it now. That dude explained it really clearly.

OpenStudy (anonymous):

yup thanks